OpenStudy (anonymous):

Solve: |x/3 + 1| = 6 A. {-7, 15} B. {-21, 15} C. {7, 15} D. {-15, 21}

OpenStudy (nurali):

|x/3 + 1| = 6 x/3+1=6 or x/3+1=-6 x/3=5 or x/3=-7 x=15 or x=-21 (15,-21)

OpenStudy (anonymous):

Thanks again :) Lol I'm so bad at this stuff

OpenStudy (nurali):

Anytime.